# Secure Control of Connected and Automated Vehicles Using Trust-Aware Robust Event-Triggered Control Barrier Functions

### 05 Jan 2024 v2

@@ Abstract
We address the security of a network of Connected and Automated Vehicles (CAVs) cooperating to safely navigate through a conflict area (e.g., traffic intersections, merging roadways, roundabouts). Previous studies have shown that such a network can be targeted by adversarial attacks causing traffic jams or safety violations ending in collisions. We focus on attacks targeting the V2X communication network used to share vehicle data and consider as well uncertainties due to noise in sensor measurements and communication channels. To combat these, motivated by recent work on the safe control of CAVs, we propose a trust-aware robust event-triggered decentralized control and coordination framework that can provably guarantee safety. We maintain a trust metric for each vehicle in the network computed based on their behavior and used to balance the tradeoff between conservativeness (when deeming every vehicle as untrustworthy) and guaranteed safety and security. It is important to highlight that our framework is invariant to the specific choice of the trust framework. Based on this framework, we propose an attack detection and mitigation scheme which has twofold benefits: (i) the trust framework is immune to false positives, and (ii) it provably guarantees safety against false positive cases. We use extensive simulations (in SUMO and CARLA) to validate the theoretical guarantees and demonstrate the efficacy of our proposed scheme to detect and mitigate adversarial attacks.
@@ I Introduction
The emergence of Connected and Automated Vehicles (CAVs) and advancements in traffic infrastructure [1] promise to offer solutions to transportation issues like accidents, congestion, energy consumption, and pollution [2], [3], [4]. To achieve these benefits, secure and efficient traffic management is crucial, particularly at bottleneck locations such as intersections, roundabouts, and merging roadways [6].
@@
We focus on decentralized algorithms as they provide manifold benefits, including added security as an attacker can only target a limited number of agents; in contrast, in a centralized scheme an attack on the central entity can potentially compromise every agent/CAV. Security of Autonomous Vehicles (AVs) has been extensively studied in existing literature [7, 8, 9] whereby the attacks can be broadly categorized into in-vehicle network attacks and V2V or V2X communication network attacks. There has been significant research done [36, 11, 12] from a control point of view with the aim of designing efficient real-time controllers for CAVs. However, ensuring security in the implementation of these controllers has received little attention, with the literature mostly limited to the security of Cooperative Adaptive Cruise Control (CACC) [17, 13, 14, 15, 19, 16]. These studies do not extend to the more critical parts of a traffic network such as intersections or roundabouts, where the repercussions of an attack are more severe, yet the literature addressing security in these cases is limited. The authors in [21] propose a technique based on public key cryptography, while [22] assesses cybersecurity risks on cooperative ramp merging by targeting V2I communication with road-side units (RSU). More comprehensive studies of the security of decentralized control and coordination algorithms for CAVs can be found in [23, 20]. In [20], an attack resilient control and coordination algorithm has been proposed using Control Barrier Functions (CBFs) without any mitigation technique. Moreover, the framework in [20] only uses V2X communication without local perception, which we deem highly useful for added security. It is also not robust to uncertainties in state estimates/measurements, which poses a security limitation as many stealthy attacks are designed to go through a Bad Data Detector (BDD) undetected.
@@
An idea that has been extensively applied to multi-agent systems is the notion of trust/reputation [27], [28], [30]. In [31] a novel trust-based CBF framework is proposed for multi-robot systems (MRSs) to provide safe control against adversarial agents; however, this cannot be directly applied to a traffic network as it is limited to a specific characterization of agents that does not apply to a road network. The authors in [32] used a trust framework to address the security of CACC. Lastly, [33] used a trust framework based on a macroscopic model of the network to tackle Sybil attacks for traffic intersections; however, the proposed framework is limited in that it provides no analysis on the fidelity of the model used for estimating traffic density to detect fake CAVs and no guarantees on preventing false positives (i.e., detecting all fake vehicles accurately and not detecting any real vehicle as fake).
The main contributions of this paper are summarized below:
@@
    We propose a novel robust trust-aware event-triggered control and coordination framework that guarantees safe coordination for CAVs in conflict areas in the presence of adversarial attacks. Our proposed formulation is robust against stealthy attacks that can pass through BDDs undetected. The benefit of event-triggered control lies in reducing the communication load, thus improving robustness against attacks.
@@
The paper is organized in six sections. The next section provides some background, followed by the threat model in Section III. In Section IV, we present the robust event-triggered control and coordination framework, which is followed by the attack mitigation in Section V. We present simulation results in Section VI. Finally, the conclusion is included in Section VII.
@@ II Background
where $x_{i}(t)$ is the distance along the lane from the origin at which CAV $i$ arrives, $v_{i}(t)$ and $u_{i}(t)$ denote the velocity and control input (acceleration/deceleration) of CAV $i$, respectively, $v_{max}>0$ denotes the maximum speed and $u_{min}<0$ is the minimum control allowed in the CZ.
A road-side unit (RSU) acts as a coordinator which receives and stores the state and control information $[x_{i}(t),v_{i}(t),u_{i}(t)]^{T}$ from CAVs through vehicle-to-infrastructure (V2X) communication. It is assumed that the coordinator knows the entry and exit lanes for each CAV upon their arrival and uses it to determine the list of MPs in its planned trajectory. It facilitates safe coordination by providing each CAV with relevant information about other CAVs in the network, particularly those that are at risk of collision.
@@ II-A Constraints/rules in the Control Zone
where $\varphi$ denotes the reaction time and $\Delta$ is a given minimum safe distance which depends on the length of these two CAVs.
@@
where $v_{min}\geq 0$ denote the minimum speed, and $u_{max}>0$ denote the maximum control allowed in the CZ respectively. $v_{max}$ and $u_{min}$ are as defined before. The coordinator finds CAV $i_{p}$ and CAV $i_{m}$ for each CAV $i\in S(t)$ from their trajectory and communicates it to CAV $i$. The determination depends on the policy adopted for sequencing CAVs whose relative performance has been studied in [35]. A common sequencing scheme is the First In First Out (FIFO) policy whereby CAVs exit the CZ in the order they arrive.
@@ II-B Decentralized control formulation:
Under this formulation, each CAV $i$ determines its control policy in a decentralized manner based on some objective that includes minimizing travel time and energy consumption, maximizing comfort, etc., governed by the dynamics (1). Expressing energy through $\frac{1}{2}u_{i}^{2}(t)$ we use $\alpha\in[0,1]$ as a relative weight between the time and energy objectives, which can be properly normalized by setting $\beta:=\frac{\alpha\max\{u_{max}^{2},u_{min}^{2}\}}{2(1-\alpha)}$ to penalize travel time relative to the energy cost of CAV $i$. Then, we can formulate an Optimal Control Problem (OCP) as follows:
@@ II-C Trust framework
$R_{i}(t)=$ $\gamma R_{i}(t-1)+\underbrace{\sum_{j\in\mathcal{B}\backslash\mathcal{B}_{a}}r_{i,j}(t)+\sum_{j\in\mathcal{B}_{a}}\prod_{k\in S_{i,j}}\tau_{k}(t)r_{i,j}(t)}_{r_{i}(t)}$
$P_{i}(t)=$ $\gamma P_{i}(t-1)+\underbrace{\sum_{j\in\mathcal{B}\backslash\mathcal{B}_{a}}p_{i,j}(t)+\sum_{j\in\mathcal{B}_{a}}\prod_{k\in S_{i,j}}\tau_{k}(t)p_{i,j}(t)}_{p_{i}(t)}\ \ \forall i\in S(t),\forall t\in[t_{i}^{0},t_{i}^{f}]$ (8)
Finally, we define a lower trust threshold $\delta\in(0,1/2)$, and a higher trust threshold $1-\delta$ for subsequent sections. It is important to emphasize that, in practice, the magnitude of negative evidence is different and significantly higher compared to the magnitude of positive evidence.
@@ III Threat model
Sybil attack A single malicious client (could be a CAV or attacker nearby the $\mathrm{CZ}$) may spoof one or multiple unique identities and register them in the coordinator queue table as detailed in [20]. Let $S_{x}(t)$ and $S_{s}(t)$ be the set of the indices of normal and fake CAVs in the FIFO queue of the coordinator unit. Therefore at any time $t$, there are $N(t)=\left|S_{x}(t)\right|+\left|S_{s}(t)\right|$ CAVs which communicate their state and control information to the coordinator. A Sybil attack is one where the $S_{s}(t)\subset S(t)$ is a nonempty set that is located in the coordinator queue table, but unknown to the coordinator.
@@
Specifically, we consider bias injection attacks as defined below:
Bias Injection (BI) attack An adversarial agent may attempt to violate safe coordination amongst CAVs, or, affect the traffic through by targetting one or more CAVs using Main-In-The-Middle (MiTM) attack by adding bias to the data sent by the CAVs to the RSU, or the data sent by the RSU to the CAVs containing state information of the relevant CAVs, or both of them. Let $\boldsymbol{y_{i}}\left(t\right)$, $i\in S(t)$ be the data (of CAV $i$, or data for CAV $i$ containing the information of the relevant CAVs) injected by the adversary during the attack; and $\boldsymbol{z_{i}}\left(t\right)$ be the actual data (of CAV $i$ sent to the RSU or data for CAV $i$ containing the information of the relevant CAVs sent by the RSU). Then, during the BI attack, $\boldsymbol{y_{i}}(t)=\boldsymbol{z_{i}}(t)+\boldsymbol{g_{i}(t)}$ where $\|\boldsymbol{g_{i}}(t)\|_{\infty}\leq\epsilon_{1}$ is the mapping used by the adversary to generate false data being stealthy.
@@ IV-A Trust-aware coordination
The trust metric is incorporated to the selected passing sequence to identify the CAVs any given CAV has to cooperate within the CZ. The cooperation with a CAV involves either constraint (2), or (3). According to this method, for every CAV $i\in S(t)$ and for every MP $j\in m_{i}$, the coordinator identifies the indices of all CAVs that precede CAV $i$ at $j$ based on the selected passing sequence until the first CAV whose trust value is greater than or equal to $1-\delta$. This leads to a new set $S_{i,j}(t)\subset S(t)$ containing all the CAV indices identified during the search process. The coordinator follows the same search process for every MP in $m_{i}$ corresponding to (3). Therefore, for each CAV $i$, the coordinator identifies $S_{i}^{p}(t)\subset S(t)$, and $S_{i}^{M}(t)=\cup_{j\in m}S_{i,j}(t)$ (where $S_{i}^{p}(t)$ is the set for (2) and $S_{i}^{M}(t)$ correspond to the set of indices for every MP) and the information is communicated to the CAV. For the example in Fig. 1, note that for CAV 4 we have $i_{p}=3$, however since $\tau_{3}<1-\delta$, the search process will continue and return $S_{4,p}=\{1,3\}$.
@@
The OCBF Controller. This approach uses the OCP formulation in (6) with each state constraint $b_{q}(\boldsymbol{x}(t))\geq 0$ mapped onto a new constraint which has the property that it is linear in the control input. Each function $b_{q}(\boldsymbol{x}(t))$ is a Control Barrier Function (CBF) derived so as to ensure the constraints (2), (3), (4) and (5) subject to the vehicle dynamics in (1) by defining $f(\boldsymbol{x}_{i}(t))=[v_{i}(t),0]^{T}$ and $g(\boldsymbol{x}_{i}(t))=[0,1]^{T}$. Each of these constraints can be easily written in the form of $b_{q}(\boldsymbol{x}_{i,j}(t))\geq 0$, $q\in\{1,2,3,4\}$ where $n$ stands for the number of constraints only dependent on state variables $\boldsymbol{x}_{i,j}(t)=[\boldsymbol{x}_{i}(t),\boldsymbol{x}_{j}(t)]^{T}$. The general form of the transformed CBF-based constraints is:
@@
where $L_{f},L_{g}$ are the Lie derivatives of a function along the system dynamics defined by $f,g$ above and $\kappa_{q}$ is a class $\mathcal{K}$ function (see [12]). By combining the OCP formulation in (6) with the CBF-based constraints of the form (10) instead of the original ones, we obtain the Optimal control with CBFs (termed OCBF) approach detailed in [38].
@@ IV-C Robust Trust-Aware CBFs
$L_{f}b_{q}(\boldsymbol{\hat{x}}_{i,j}(t)-\boldsymbol{w}_{i,j}(t))+L_{g}b_{q}(\boldsymbol{\hat{x}}_{i,j}(t)-\boldsymbol{w}_{i,j}(t))u_{i}(t)+\kappa_{q,\tau_{j}}(b_{q}(\boldsymbol{\hat{x}}_{i,j}(t)-\boldsymbol{w}_{i,j}(t)))\geq 0.$ (13)
@@
$\hat{v}_{i_{p}}(t)+w_{i_{p}}^{(v)}(t)-\hat{v}_{i}(t)-w_{i}^{(v)}(t)-\varphi u_{i}(t)-\kappa_{q,\tau_{i_{p}}}(\hat{x}_{i_{p}}(t)+w_{i_{p}}^{(x)}(t)-\hat{x}_{i}(t)-w_{i}^{(x)}(t)-\varphi\hat{v}_{i}(t)-\varphi w_{i}^{(v)}(t)-\Delta)\geq 0$
@@
$\min_{\{\boldsymbol{w}_{i,j}(t):\|\boldsymbol{w}_{i,j}(t)\|_{\infty}\leq\epsilon_{1}\}}[L_{f}b_{q}(\boldsymbol{\hat{x}_{i,j}}(t)-\boldsymbol{w}_{i,j}(t))]+L_{g}b_{q}(\boldsymbol{\hat{x}_{i,j}}(t)-\boldsymbol{w}_{i,j}(t))u_{i}(t)+\kappa_{q,\tau_{j}}(b_{q}(\boldsymbol{\hat{x}_{i,j}}(t)-\boldsymbol{w}_{i,j}(t)))]\geq 0$ (15)
@@
The satisfaction of (15) guarantees the satisfaction of the constraint (13) (and (10)) since it is a lower bound for (10) which according to Theorem 1 in [12] makes the set $C$ forward invariant $\forall t\geq t_{0}$ w.r.t (1). ∎

### 25 Mar 2024 v3

@@ Abstract
We address the security of a network of Connected and Automated Vehicles (CAVs) cooperating to safely navigate through a conflict area (e.g., traffic intersections, merging roadways, roundabouts). Previous studies have shown that such a network can be targeted by adversarial attacks causing traffic jams or safety violations ending in collisions. We focus on attacks targeting the V2X communication network used to share vehicle data and consider as well uncertainties due to noise in sensor measurements and communication channels. To combat these, motivated by recent work on the safe control of CAVs, we propose a trust-aware robust event-triggered decentralized control and coordination framework that can provably guarantee safety. We maintain a trust metric for each vehicle in the network computed based on their behavior and used to balance the tradeoff between conservativeness (when deeming every vehicle as untrustworthy) and guaranteed safety and security. It is important to highlight that our framework is invariant to the specific choice of the trust framework. Based on this framework, we propose an attack detection and mitigation scheme which has twofold benefits: (i) the trust framework is immune to false positives, and (ii) it provably guarantees safety against false positive cases. We use extensive simulations (in SUMO and CARLA) to validate the theoretical guarantees and demonstrate the efficacy of our proposed scheme to detect and mitigate adversarial attacks. The code for the simulated scenarios can be found in this link.
publicationid: pubid: Symposium on Vehicles Security and Privacy (VehicleSec) 2024 26 February 2024, San Diego, CA, USA ISBN 979-8-9894372-7-6 https://dx.doi.org/10.14722/vehiclesec.2024.23037 www.ndss-symposium.org
@@ I Introduction
The emergence of Connected and Automated Vehicles (CAVs) and advancements in traffic infrastructure [1] promise to offer solutions to transportation issues like accidents, congestion, energy consumption, and pollution [2, 3]. To achieve these benefits, secure and efficient traffic management is crucial, particularly at bottleneck locations such as intersections, roundabouts, and merging roadways [4].
@@
We focus on decentralized algorithms as they provide manifold benefits, including added security since an attacker can only target a limited number of agents; in contrast, in a centralized scheme an attack on the central entity can potentially compromise every agent/CAV. Security of Autonomous Vehicles (AVs) has been extensively studied in existing literature [5, 6, 7] whereby the attacks can be broadly categorized into in-vehicle network attacks and V2V or V2X communication network attacks. There has been significant research done [8, 9, 10] from a control point of view with the aim of designing efficient real-time controllers for CAVs. However, ensuring security in the implementation of these controllers has received little attention, with the literature mostly limited to the security of Cooperative Adaptive Cruise Control (CACC) [11, 12, 13, 14, 15]. These studies do not extend to the more critical parts of a traffic network such as intersections or roundabouts, where the repercussions of an attack are more severe, yet the literature addressing security in these cases is limited.
The authors in [16] propose a technique based on public key cryptography, while [17] assesses cybersecurity risks on cooperative ramp merging by targeting V2I communication with road-side units (RSU). More comprehensive studies of the security of decentralized control and coordination algorithms for CAVs can be found in [18, 19]. In [19], an attack resilient control and coordination algorithm is proposed using Control Barrier Functions (CBFs) without any mitigation technique. Moreover, the framework in [19] only uses V2X communication without local perception, which we deem highly useful for added security. It is also not robust to uncertainties in state estimates/measurements, which poses a security limitation as many stealthy attacks are designed to evade detection by a Bad Data Detector (BDD).
@@
The notion of trust/reputation has been applied to multi-agent systems including Intelligent Transportation Systems (ITS) in [20, 21, 22, 23]. In [24] a novel trust-based CBF framework is proposed for multi-robot systems (MRSs) to provide safe control against adversarial agents. However, this cannot be directly applied to a traffic network as it is limited to a specific characterization of agents that does not apply to a road network. The authors in [25] propose a trust framework to address the security of CACC. Along a similar vein, [26] employs a trust framework to address Sybil attacks within traffic intersections using a macroscopic network model. However, it is constrained by the accuracy of the traffic density estimation model in detecting fake vehicles (CAVs) and also offers no guarantees on preventing false positives (i.e., detecting all fake vehicles accurately and not detecting any real vehicle as fake).
In our paper we tackle the aforementioned shortcomings beginning with making the control and coordination robust against disturbances and uncertainties in the states measurements and estimations. Besides that, we also incorporate mitigation in order to subside the effect of the attacks on the network performance. Thus our main contributions are summarized below:
@@
    We propose a novel robust trust-aware event-triggered control and coordination framework that guarantees safe coordination for CAVs in conflict areas in the presence of adversarial attacks.
@@
    Our proposed formulation is robust against stealthy attacks that can pass through BDDs undetected. The benefit of event-triggered control lies in reducing the communication load, thus improving robustness against attacks.
  3. 3.
@@
While our framework views security as a specification in a control and coordination problem, it is important to note that various network security measures like cryptographic techniques can complement this framework. The paper is organized in six sections. The next section provides some background, followed by the threat model in Section III. In Section IV, we present the robust event-triggered control and coordination framework, which is followed by the attack mitigation in Section V. We present simulation results in Section VI. Finally, the conclusion is included in Section VII.
@@ II Background
where $x_{i}(t)$ is the distance along the lane from the origin at which CAV $i$ arrives, $v_{i}(t)$ and $u_{i}(t)$ denote the velocity and control input (acceleration/deceleration) of CAV $i$, respectively.
A road-side unit (RSU) acts as a coordinator which receives and stores the state and control information $[x_{i}(t),v_{i}(t),u_{i}(t)]^{T}$ from CAVs through vehicle-to-infrastructure (V2X) communication. Additionally, it also stores and updates the a trust metric for each CAV in the CZ. It is assumed that the coordinator knows the entry and exit lanes for each CAV upon their arrival and uses it to determine the list of MPs in its planned trajectory. It facilitates safe coordination by providing each CAV with relevant information about other CAVs in the network, particularly those that are at risk of collision.
@@ II-A Constraints/rules in the Control Zone
where $\varphi$ denotes the reaction time and $\Delta\in\mathbb{R}_{>0}$ is a given minimum safe distance which depends on the length of these two CAVs.
@@
where $v_{min}\geq 0$ and $v_{max}>0$ denote the minimum and maximum allowed speed in the CZ respectively, and $u_{min}<0$ $u_{max}>0$ denote the minimum and maximum de(ac)celeration allowed in the CZ respectively. The coordinator finds CAV $i_{p}$ and CAV $i_{m}$ for each CAV $i\in S(t)$ from their trajectory and communicates it to CAV $i$. The determination depends on the policy adopted for sequencing CAVs whose relative performance has been studied in [28]. A common sequencing scheme is the First In First Out (FIFO) policy whereby CAVs exit the CZ in the order they arrive.
@@ II-B Decentralized control formulation:
Under this formulation, each CAV $i$ determines its control policy in a decentralized manner based on some objective that includes minimizing travel time and energy consumption, maximizing comfort, etc., governed by the dynamics (1). Expressing energy through $\frac{1}{2}u_{i}^{2}(t)$ we use $\alpha\in[0,1]$ as a relative weight between the time and energy objectives, which can be properly normalized by setting $\beta:=\frac{(1-\alpha)\max\{u_{max}^{2},u_{min}^{2}\}}{2\alpha}$ to penalize travel time relative to the energy cost of CAV $i$. Then, we can formulate an Optimal Control Problem (OCP) as follows:
@@ II-C Trust framework
$R_{i}(t)=$ $\gamma R_{i}(t-1)+\sum_{j\in\mathcal{B}\backslash\mathcal{B}_{a}}r_{i,j}(t)+\sum_{j\in\mathcal{B}_{a}}\prod_{k\in S_{i,j}}\tau_{k}(t)r_{i,j}(t)$
$P_{i}(t)=$ $\gamma P_{i}(t-1)+\underbrace{\sum_{j\in\mathcal{B}\backslash\mathcal{B}_{a}}p_{i,j}(t)+\sum_{j\in\mathcal{B}_{a}}\prod_{k\in S_{i,j}}\tau_{k}(t)p_{i,j}(t)}_{p_{i}(t)}$
$\forall i\in S(t),\forall t\in[t_{i}^{0},t_{i}^{f}]$ (8)
Finally, we define a lower trust threshold $\delta\in(0,1/2)$, and a higher trust threshold $1-\delta$ for subsequent sections. It is important to emphasize that, in practice, the magnitude of negative evidence is different and significantly higher compared to the magnitude of positive evidence. Note that, every vehicle is deemed untrustworthy i.e. $\tau_{i}(t_{i}^{0})=0\ \forall i$ upon arrival in the CZ. The coordinator updates the trust for every CAV based on the outcome of the behavior specification checks. The zero trust model is used to prioritize safety in our proposed framework.
@@ III Threat model
Sybil attack A single adversarial agent (could be a CAV or attacker nearby the $\mathrm{CZ}$) may spoof one or multiple unique identities and register them in the coordinator queue table as detailed in [19]. Let $S_{x}(t)$ and $S_{s}(t)$ be the set of the indices of normal and fake CAVs in the FIFO queue of the coordinator unit. Therefore at any time $t$, there are $N(t)=\left|S_{x}(t)\right|+\left|S_{s}(t)\right|$ CAVs which communicate their state and control information to the coordinator. A Sybil attack is one where the $S_{s}(t)\subset S(t)$ is a nonempty set that is located in the coordinator queue table, but unknown to the coordinator.
@@
Specifically, we consider bias injection attacks as defined below:
Bias Injection (BI) attack An adversarial agent may attempt to violate safe coordination amongst CAVs, or affect the traffic by targeting one or more CAVs using Person-In-The-Middle attack by adding bias to the data sent by the CAVs to the RSU, or the data sent by the RSU to the CAVs containing state information of the relevant CAVs, or both of them. Let $\boldsymbol{y_{i}}\left(t\right)$, $i\in S(t)$ be the data (of CAV $i$, or data for CAV $i$ containing the information of the relevant CAVs) injected by the adversary during the attack; and $\boldsymbol{z_{i}}\left(t\right)$ be the actual data (of CAV $i$ sent to the RSU or data for CAV $i$ containing the information of the relevant CAVs sent by the RSU). Then, during the BI attack, $\boldsymbol{y_{i}}(t)=\boldsymbol{z_{i}}(t)+\boldsymbol{g_{i}(t)}$ where $\|\boldsymbol{g_{i}}(t)\|_{\infty}\leq\epsilon_{1}$ is the mapping used by the adversary to generate false data being stealthy.
@@ IV-A Trust-aware coordination
The coordinator computes and updates the trust metric for each CAV in the CZ as shown in figure 1. The trust metric is incorporated to the selected passing sequence to identify the CAVs any given CAV has to cooperate within the CZ. The cooperation with a CAV involves either constraint (2), or (3). According to this method, for every CAV $i\in S(t)$ and for every MP $j\in m_{i}$, the coordinator identifies the indices of all CAVs that precede CAV $i$ at $j$ based on the selected passing sequence until the first CAV whose trust value is greater than or equal to $1-\delta$. This leads to a new set $S_{i,j}(t)\subset S(t)$ containing all the CAV indices identified during the search process. The coordinator follows the same search process for every MP in $m_{i}$ corresponding to (3). Therefore, for each CAV $i$, the coordinator identifies $S_{i}^{p}(t)\subset S(t)$, and $S_{i}^{M}(t)=\cup_{j\in m}S_{i,j}(t)$ (where $S_{i}^{p}(t)$ is the set for (2) and $S_{i}^{M}(t)$ correspond to the set of indices for every MP) and the information is communicated to the CAV. For the example in Fig. 1, note that for CAV 4 we have $i_{p}=3$, however since $\tau_{3}<1-\delta$, the search process will continue and return $S_{4,p}=\{1,3\}$.
@@
The OCBF Controller. This approach uses the OCP formulation in (6) with each state constraint $b_{q}(\boldsymbol{x}(t))\geq 0$ mapped onto a new constraint which has the property that it implies $b_{q}(\boldsymbol{x}(t))\geq 0$ and it is linear in the control input. The function $b_{q}(\boldsymbol{x}(t))$ is called a Control Barrier Function (CBF) [30]. We use such CBFs so as to ensure the constraints (2), (3), (4) and (5) are satisfied subject to the vehicle dynamics in (1) by defining $f(\boldsymbol{x}_{i}(t))=[v_{i}(t),0]^{T}$ and $g(\boldsymbol{x}_{i}(t))=[0,1]^{T}$. Each of these constraints can be easily written in the form of $b_{q}(\boldsymbol{x}_{i,j}(t))\geq 0$, $q\in\{1,2,3,4\}$ where $n$ stands for the number of constraints only dependent on state variables $\boldsymbol{x}_{i,j}(t)=[\boldsymbol{x}_{i}(t),\boldsymbol{x}_{j}(t)]^{T}$. The general form of the transformed CBF-based constraints is:
@@
where $L_{f},L_{g}$ are the Lie derivatives of a function along the system dynamics defined by $f,g$ above and $\kappa_{q}$ is a class $\mathcal{K}$ function. By combining the OCP formulation in (6) with the CBF-based constraints of the form (10) instead of the original ones, we obtain the Optimal control with CBFs (termed OCBF) approach detailed in [10].
@@ IV-C Robust Trust-Aware CBFs
$L_{f}b_{q}(\boldsymbol{\hat{x}}_{i,j}(t)-\boldsymbol{w}_{i,j}(t))+L_{g}b_{q}(\boldsymbol{\hat{x}}_{i,j}(t)-\boldsymbol{w}_{i,j}(t))u_{i}(t)+$
$\kappa_{q,\tau_{j}}(b_{q}(\boldsymbol{\hat{x}}_{i,j}(t)-\boldsymbol{w}_{i,j}(t)))\geq 0.$ (13)
@@
$\hat{v}_{i_{p}}(t)+w_{i_{p}}^{(v)}(t)-\hat{v}_{i}(t)-w_{i}^{(v)}(t)-\varphi u_{i}(t)-\kappa_{q,\tau_{i_{p}}}(\hat{x}_{i_{p}}(t)$
$+w_{i_{p}}^{(x)}(t)-\hat{x}_{i}(t)-w_{i}^{(x)}(t)-\varphi\hat{v}_{i}(t)-\varphi w_{i}^{(v)}(t)-\Delta)\geq 0$
@@
$\min_{\{\boldsymbol{w}_{i,j}(t):\|\boldsymbol{w}_{i,j}(t)\|_{\infty}\leq\epsilon_{1}\}}[L_{f}b_{q}(\boldsymbol{\hat{x}_{i,j}}(t)-\boldsymbol{w}_{i,j}(t))]+L_{g}b_{q}(\boldsymbol{\hat{x}_{i,j}}(t)$ (15)
$-\boldsymbol{w}_{i,j}(t))u_{i}(t)+\kappa_{q,\tau_{j}}(b_{q}(\boldsymbol{\hat{x}_{i,j}}(t)-\boldsymbol{w}_{i,j}(t)))]\geq 0$
@@
The satisfaction of (15) guarantees the satisfaction of the constraint (IV-C) (and (10)) since it is a lower bound for (10) which according to Theorem 1 in [10] makes the set $C$ forward invariant $\forall t\geq t_{0}$ w.r.t (1). ∎
Based on the information in the table (as shown in Fig. 1) the coordinator communicates the state information and the trust value of the CAVs in $S_{i}^{p}$ and $S_{i}^{M}$ corresponding to constraints (2) and (3) respectively to each CAV $i$ in the CZ.